Find the length of a rectangular lot with a perimeter of 118m if the length is 5m more than the width.
Answer by oberobic(2304)

Perimeter = 2*(Length + Width)
.
L = W + 5
.
P = 118 m
.
2*(L+W) = 118
.
L+W = 59
.
Substitute L = W+5
.
W+5 +W = 59
.
2W = 54
.
W = 27
.
L = W+5 = 27+5 = 32
.
Answer: Length equals 32.
.
Done.